# Harnessing Wavelet Transformations for Generalizable Deepfake Forgery Detection

### v2

@@ Abstract
Index Terms—  Face Forgery, Deepfakes, Self-supervision, Wavelet Transforms, Contastive-Language Image Pretraining (CLIP).
@@ 1 Introduction
In today’s digital landscape, we are witnessing an inundation of counterfeit images, arising from various sources. Some of these images are manipulated versions of authentic photos, altered using tools such as FaseShifter [1] and proprietor photoshop tools [2], while others are crafted through advanced machine learning algorithms. The advent and refinement of deep generative models [3, 4, 5] have particularly highlighted the latter category, drawing both admiration for the photo-realistic images they can produce and concern over their potential misuse. The challenge is compounded by the diverse origins of these fake images; they may manifest as real human faces created by generative adversarial networks or as intricate scenes synthesized by diffusion models [5]. This growing variety underscores the inevitability of encountering new forms of image forgery. Against this challenges, our research aims to devise a diverse generalizable fake detection framework capable of identifying any falsified image, even when training is confined to a single type of generative model.
@@
There are numerous methods developed for deepfake generalization both within and cross-domain evaluation [6, 7, 8, 9, 10]. Of which, some works rely of basic encoders such as EfficentNet [9] and Xception [11]. Some models leverage frequency-based statistics for identfying some details which spatial domain can’t capture [12, 13, 14]. Most of the existing deepfake detection models demonstrate significant results in scenarios where the training and testing data come from the same dataset. However, these detectors frequently face challenges in cross-domain or cross-dataset scenarios, where there is a significant discrepancy between the distribution of the training data and that of the testing data. Our method significantly advances the field of digital forensics by offering a robust model capable of countering the evolving threat of digital image forgery. It achieves this by effectively generalizing across different datasets and adeptly identifying deepfakes produced by powerful, previously unseen generators.
The proposed approach offers distinct advantages over current methodologies in two key directions:
@@
Fig. 1: Wavelet-CLIP: The comprehensive workflow of the proposed Wavelet-CLIP. Initially, the model ingests real and counterfeit image samples, which are processed by a ViT-L/14 encoder, pretrained with CLIP weights [15], to produce feature representations. These representations are then subjected to Discrete Wavelet Transform (DWT) to downsample into low-frequency and high-frequency components. The low-frequency component is further refined using a MLP keeping the high frequency features $fv_{\text{high}}$ constant (where, the ”$=$” signifies an identity mapping). Subsequently, the transformed representations are processed by another MLP to classify the image is a deepfake or genuine.
@@ 2.1 Encoder
A good encoder has to understand the crucial features from the image distribution and map them to the latent space. This latent features should carry the prominent features of image. But, when it comes to generalization, the features have to be more relevant irrespective of trained or seen samples. In such scenarios, model that is trained on internet-scale data in a self-supervised fashion should provide fine-grained features irrespective of nature of data. Hence, we adopt pre-trained vision transformer [16] model that is trained via. CLIP fashion[15] pertaining strong one-shot transferable features. This encoder maps a image into a representation space of feature dimension $d$ where, $Enc_{\phi}:\mathbb{R}^{256\times 256\times 3}\to\mathbb{R}^{d}$ (we denote frozen encoder as $\phi$). The latent features $Z$ captured for‘our study are using ViT-L/14 [15] and is represented as,
@@ 2.2 Classification Head
CLIP [20] CVPR-23 ViT [16] Self-Supervised 0.743 0.750 0.730 0.747
@@
Algorithm 1 Wavelet-CLIP
1:  Input: Dataset $\mathcal{D}$, trained $Enc_{\phi}^{(\text{ViT})}(.)$, $\epsilon$
2:  for $n=1$ to $\epsilon$ do
3:     $Z^{(n)}=Enc_{\phi}^{(\text{ViT})}(x^{(n)})$
4:     $fv_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}=\text{DWT}(Z^{(n)})$
5:     ${fv^{\prime}}_{\text{low}}^{(n)}=\text{MLP}(fv_{\text{low}}^{(n)})$
6:     $Z_{\text{new}}^{(n)}=\text{IDWT}([{fv^{\prime}}_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}])$
7:     $\text{cls}_{n}=\text{MLP}(Z_{\text{new}}^{(n)})$
8:  end for
9:  return $\text{cls}_{n}$
Wavelet Transforms
Wavelet Transforms are used to analyze various frequency components of a signal and is particularly useful representations that have hierarchical or multi-scale structure [27]. Applying a Discrete Wavelet Transform (DWT) the representation splits into low and high frequency components. Low-frequency components are responsible in capturing broad and nuanced features. Whereas high frequency components capture sharp features. The DWT decomposes as Low-pass filter ($l$), that captures the low-frequency component. Similarly, captures the high-frequency component, resulting in a down-sampled signal $d_{1}$. These are denoted as,
$s_{1k}$ $=\sum_{j}l_{(j-2k)}s_{j}$ (2)
$d_{1k}$ $=\sum_{j}h_{(j-2k)}s_{j}$ (3)
These operations can be expressed in matrix form, where $L$ and $H$ are matrices constructed from low-pass and high-pass filter coefficients, respectively, performing filtering and down-sampling in a single step through the operations $s_{1}=Ls$ and $d_{1}=Hs$. The Inverse Discrete Wavelet Transform IDWT reconstructs the original signal $s$ from its low-frequency ($s_{1}$) and high-frequency ($d_{1}$) components using:
$s_{j}=\sum_{k}(l_{j-2k}s_{1k}+h_{j-2k}d_{1k})$ (4)
For 2D signals like images, applying DWT along both dimensions results in four components–$X_{ll},$ (low-frequency in both dimensions), $X_{lh}$ (low-frequency row-wise, high-frequency column-wise), $X_{hl}$ (high-frequency row-wise, low-frequency column-wise), and $X_{hh}$ (high-frequency in both dimensions)—by applying matrices $L$ and $H$ across rows and columns, respectively. The DWT and IDWT can be expressed as,
$X_{ll}$ $=LXL^{T},$ (5)
$X_{lh}$ $=HXL^{T},$ (6)
$X_{hl}$ $=LXH^{T},$ (7)
$X_{hh}$ $=HXH^{T},$ (8)
$X=L^{T}X_{ll}L+H^{T}X_{lh}L+L^{T}X_{hl}H+H^{T}X_{hh}H.$ (9)
@@
Wavelet Classifier
$fv_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}$ $=\text{DWT}(Z^{(n)})$ (10)
${fv^{\prime}}_{\text{low}}^{(n)}$ $=\text{MLP}(fv_{\text{low}}^{(n)})$ (11)
3 Experiments
In this section, we will first detail the evaluation protocol, datasets used. Later we will discuss the performance of our approach with various state-of-the-art approaches.
In alignment with the training and evaluation protocol established by Yan et al. [6], the models undergo initial training on the FaceForensics++ c23 dataset [19]. Subsequent evaluations employ a cross-domain test using datasets such as Celeb-DF v1 (CDFv1) [28], Celeb-DF v2 (CDFv2) [28], and FaceShifter (Fsh) [1], thereby providing a robust framework to test the generalization capabilities of all the models. While the existing benchmarks do not encompass tests on emerging diffusion models, our research extends to examining generalizability with novel synthetic samples. Utilizing state-of-the-art diffusion models like DDPM [3], DDIM [4], and LDM [5] (wihtout text-guidance), we generate 10,000 images (from each model) from the CelebA dataset–none of which were included in training phase. This approach enables a comprehensive evaluation of model’s adaptability to novel and unseen data, leveraging its potential for practical deployment in digital forensics. To assess the effectiveness of the results we use AUC (Area Under the Curve) and EER (Equal Error Rate) as fundamental metrics [6].
@@ 3 Experiments
Table 1 clearly demonstrates the superior performance of Wavelet-CLIP compared to existing models. Specifically, for the CDFv1 dataset, traditional CLIP features fall short in capturing detailed representations; a standard ViT-L model [20] achieves an AUC of 0.743, while our model shows a significant improvement with a $+1.3\%$ increase. In every other scenario, our Wavelet-CLIP model stands out by consistently delivering strong performance. Notably, transformer-based models, including ours and those developed by Ojha [20], demonstrate effective representation capturing abilities for the FaceShifter (Fsh) dataset [1]. Interestingly, the best non-transformer model, Face X-ray [25], achieves an AUC of 0.655, whereas our approach exhibits a significant improvement with a $+7.7\%$ increase in AUC. The Fsh dataset, known for its sophisticated face manipulation techniques across diverse scenarios, presents a substantial challenge; yet, it appears that pre-trained transformers are particularly adept at discerning the subtle forgeries inherent in such deepfakes.
@@
Next, Table 2 assesses the performance of Wavelet-CLIP on face images generated by unseen diffusion-based models. The aggregated performance of Wavelet-CLIP outperforms standard CLIP [20], showing an $+1.3\%$ increase in AUC and a $-1.2\%$ reduction in EER respectively. This highlights the substantial impact of integrating wavelet transformations within the classification head. Additionally, among the non-transformer based models, Capsule [17]—noted for its rotational invariance capabilities—performs best, yet it still falls short of matching Wavelet-CLIP, with performance differences of $12.5\%$ in AUC and $10.5\%$ in EER respectively. This further underscores the superior effectiveness of Wavelet-CLIP in handling sophisticated generative challenges.

### v3

# Wavelet-Driven Generalizable Framework for Deepfake Face Forgery Detection

@@ 1 Introduction
In today’s digital landscape, we are witnessing an inundation of counterfeit images, arising from various sources. Some of these images are manipulated versions of authentic photos, altered using tools such as FaseShifter [9] and proprietor photoshop tools [15], while others are crafted through advanced machine learning algorithms. The advent and refinement of deep generative models [20, 14, 13] have particularly highlighted the latter category, drawing both admiration for the photo-realistic images they can produce and concern over their potential misuse. The challenge is compounded by the diverse origins of these fake images; they may manifest as real human faces created by generative adversarial networks or as intricate scenes synthesized by diffusion models [13]. This growing variety underscores the inevitability of encountering new forms of image forgery.
The proliferation of diffusion models [20, 14, 13] has revolutionized the field of generative AI, enabling the creation of highly realistic synthetic images with exceptional quality and diversity. These models have demonstrated remarkable capabilities in producing photorealistic human faces, complex natural scenes, and seamlessly manipulated content. However, this rapid advancement has also raised significant concerns regarding their potential misuse for malicious purposes, such as generating deepfake media or spreading disinformation. As diffusion models continue to evolve, it becomes increasingly challenging to distinguish between real and fake images, amplifying the need for robust fake image detection systems that can generalize across diverse generative families. Against this challenges, our research aims to devise a diverse generalizable fake detection framework capable of identifying any falsified image, even when training is confined to a single type of generative model.
Traditionally, fake image detection has been approached as a binary classification task, where deep neural networks are trained to distinguish real images from synthetic ones generated by a specific model, such as diffusion or GAN methods. While these approaches excel within the same generative family (e.g., detecting fake images produced by diffusion variants like LDM [13] or Guided Diffusion [6]), they fail to generalize when exposed to unseen generative families. This limitation arises because these classifiers tend to rely on low-level artifacts unique to the training model, often referred to as ”fingerprints.” Consequently, fake images generated by alternative methods that lack these specific fingerprints are misclassified as real, leading to a skewed decision boundary and poor generalization to novel image generation techniques.
@@
There are numerous methods developed for deepfake generalization both within and cross-domain evaluation [33, 27, 1, 32, 17]. Of which, some works rely of basic encoders such as EfficentNet [32] and Xception [3]. Some models leverage frequency-based statistics for identfying some details which spatial domain can’t capture [12, 8, 10]. Most of the existing deepfake detection models demonstrate significant results in scenarios where the training and testing data come from the same dataset. However, these detectors frequently face challenges in cross-domain or cross-dataset scenarios, where there is a significant discrepancy between the distribution of the training data and that of the testing data.
Our method significantly advances the field of digital forensics by offering a robust model capable of countering the evolving threat of digital image forgery. To address these limitations,we avoid explicit learning for real-vs-fake image classification and utilize the feature space of a large pre-trained vision-language model, CLIP-ViT [28] which has been trained on internet-scale datasets for tasks unrelated to fake detection. It achieves this by effectively generalizing across different datasets and adeptly identifying deepfakes produced by powerful, previously unseen generators. The proposed approach offers distinct advantages over current methodologies in two key directions:
@@
Figure 1: Wavelet-CLIP: The comprehensive workflow of the proposed Wavelet-CLIP. Initially, the model ingests real and counterfeit image samples, which are processed by a ViT-L/14 encoder, pretrained with CLIP weights [28], to produce feature representations. These representations are then subjected to Discrete Wavelet Transform (DWT) to downsample into low-frequency and high-frequency components. The low-frequency component is further refined using a MLP keeping the high frequency features $fv_{\text{high}}$ constant (where, the ”$=$” signifies an identity mapping). Subsequently, the transformed representations are processed by another MLP to classify the image is a deepfake or genuine.
2 Related Works
2.1 Naive Detectors
Naive detectors utilize the existing state-of-the-art CNN architectures to directly classify images as real or fake. These models do not rely on any handcrafted layers or domain-specific knowledge but instead learn features directly from the data during training. MesoNet [1] and MesoInception [1], which are lightweight networks optimized for efficiency and designed to capture mesoscopic features indicative of manipulations. More advanced naive detectors, such as Xception [3] and EfficientNet-B4 [32], employ modern CNN architectures that are computationally efficient.
Naive detectors are simple to implement and often achieve reasonable performance but may struggle with generalization across datasets and sophisticated forgeries due to their reliance on learned features without deeper semantic insights.
2.2 Spatial Detectors
Spatial detectors focus on analyzing the spatial domain of images, often employing advanced techniques to detect localized artifacts introduced during manipulations. Models like Capsule Networks [26] leverage dynamic routing between capsules to model spatial hierarchies irrespective of rotations, while DSP-FWA [22] specializes in detecting warping artifacts that occur during face-swapping manipulations. Face X-ray [21] zeroes in on boundary artifacts between manipulated and non-manipulated regions, leveraging high-resolution features to isolate forgery artifacts. Models like CORE [27] and UCF [17] emphasize learning robust and consistent features, enhancing generalization to unseen forgeries.
By focusing on spatial inconsistencies such as blending, texture mismatches, or altered facial regions, spatial detectors excel at identifying specific manipulations but may require more sophisticated pre-processing and training strategies. These methods might be robust to visually perceptually forgeries but, can’t perceive hidden forgeries.
2.3 Frequency Detectors
Frequency detectors analyze the frequency domain of images to detect subtle artifacts not visible in the spatial domain. These models address the limitations of spatial detectors by capturing inconsistencies in high-frequency components, noise patterns, and phase information. For instance, F3Net [12] uses adaptive filters to mine forgery clues in the frequency domain, making it effective at identifying hidden noise introduced during manipulations. SRM [10] employs high-frequency filters to identify subtle pixel-level inconsistencies.
Frequency detectors are particularly robust and operate in the frequency domain to identify manipulations by detecting artifacts in noise patterns or phase spectra. This complements spatial detectors by providing insights into overlooked forgery artifacts.
2.4 Generalizable Detectors
In Yan et al. [33] study, the methods focus on cross-domain generalization but, deepfakes can emerge from nowhere. Thus, a generalizable model should have the capability of identifying unseen fake or forgery images. To address this challenge Ojha et al. [11] provided a new direction of solving unseen deepfakes generated from diffusion and autoregressive methods. Unlike traditional classifiers trained explicitly for real-vs-fake classification, which fail to generalize to new generative model families, the proposed approach leverages feature spaces from large, pre-trained vision-language models such as CLIP [28]. Later, cozzolino et al. [4] have comprehensively analyzed the frozen CLIP features are performed exhaustive experiments on various unseen deepfakes and showcased the significance of frozen CLIP features for generalizable deepfake identification.
@@ 3.1 Encoder
A good encoder has to understand the crucial features from the image distribution and map them to the latent space. These latent features should carry the prominent features of the image. But, when it comes to generalization, the features have to be more relevant irrespective of trained or seen samples. In such scenarios, a model that is trained on internet-scale data in a self-supervised fashion should provide fine-grained features irrespective of the nature of the data. Hence, we adopt a pre-trained vision transformer [7] model that is trained via. CLIP fashion[28] pertaining strong one-shot transferable features. This encoder maps an image into a representation space of feature dimension $d$ where $Enc_{\phi}:\mathbb{R}^{256\times 256\times 3}\to\mathbb{R}^{d}$ (we denote frozen encoder as $\phi$). The latent features $Z$ captured for our study are using ViT-L/14 [28] and is represented as,
@@ 3.2 Classification Head
Algorithm 1 Wavelet-CLIP
1:  Input: Dataset $\mathcal{D}$, Encoder $Enc_{\phi}^{(\text{ViT})}(.)$, $\epsilon$, $n$;
2:  for $\textsc{iterations}=1$ to $\epsilon$ do
3:     for $\textsc{batch}=n$ do
4:        $Z^{(n)}=Enc_{\phi}^{(\text{ViT})}(x^{(n)})$
5:        $fv_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}=\text{DWT}(Z^{(n)})$
6:        ${fv^{\prime}}_{\text{low}}^{(n)}=\text{MLP}(fv_{\text{low}}^{(n)})$
7:        $Z_{\text{new}}^{(n)}=\text{IDWT}([{fv^{\prime}}_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}])$
8:        $\text{cls}_{n}=\text{MLP}(Z_{\text{new}}^{(n)})$
9:     end for
10:  end for
11:  return $\text{cls}_{n}$
Wavelet Transforms
Wavelet Transforms are used to analyze various frequency components of a signal and is particularly useful representations that have hierarchical or multi-scale structure [24]. Applying a Discrete Wavelet Transform (DWT) the representation splits into low and high frequency components. Low-frequency components are responsible in capturing broad and nuanced features. Whereas high frequency components capture sharp features.
The Discrete Wavelet Transform (DWT) of a one-dimensional signal $s=\{s_{j}\}_{j\in\mathbb{Z}}$ decomposes it into two components: a low-frequency approximation $s_{1}=\{s_{1k}\}_{k\in\mathbb{Z}}$ and a high-frequency detail $d_{1}=\{d_{1k}\}_{k\in\mathbb{Z}}$. These components are defined as,
$s_{1k}=\sum_{j\in\mathbb{Z}}l_{j-2k}s_{j},\quad d_{1k}=\sum_{j\in\mathbb{Z}}h_{j-2k}s_{j},$ (2)
where $l=\{l_{k}\}_{k\in\mathbb{Z}}$ and $h=\{h_{k}\}_{k\in\mathbb{Z}}$ represent the low-pass and high-pass filters, respectively, associated with an orthogonal wavelet. The Inverse Discrete Wavelet Transform (IDWT) allows the reconstruction of the original signal $s=\{s_{j}\}_{j\in\mathbb{Z}}$ from its low-frequency approximation $s_{1}=\{s_{1k}\}_{k\in\mathbb{Z}}$ and high-frequency detail $d_{1}=\{d_{1k}\}_{k\in\mathbb{Z}}$. The reconstruction is performed as,
$s_{j}=\sum_{k\in\mathbb{Z}}\left(l_{j-2k}s_{1k}+h_{j-2k}d_{1k}\right),$ (3)
where $l=\{l_{k}\}_{k\in\mathbb{Z}}$ and $h=\{h_{k}\}_{k\in\mathbb{Z}}$ are the low-pass and high-pass filters associated with the orthogonal wavelet. In this equation, $l_{j-2k}$ acts as an interpolation filter that reconstructs the low-frequency components,$h_{j-2k}$ reconstructs the high-frequency details, the summation runs over all integers $k$, ensuring that both low-frequency and high-frequency components contribute to the reconstructed signal $s_{j}$.
These operations can be expressed in matrix form, where $L$ and $H$ are matrices constructed from low-pass and high-pass filter coefficients, respectively. For 2D signals like images, applying DWT along both dimensions results in four components–$X_{ll},$ (low-frequency in both dimensions), $X_{lh}$ (low-frequency row-wise, high-frequency column-wise), $X_{hl}$ (high-frequency row-wise, low-frequency column-wise), and $X_{hh}$ (high-frequency in both dimensions)—by applying matrices $L$ and $H$ across rows and columns, respectively. The DWT and IDWT can be expressed as,
$X_{ll}$ $=LXL^{T},$ (4)
$X_{lh}$ $=HXL^{T},$ (5)
$X_{hl}$ $=LXH^{T},$ (6)
$X_{hh}$ $=HXH^{T},$ (7)
$X=L^{T}X_{ll}L+H^{T}X_{lh}L+L^{T}X_{hl}H+H^{T}X_{hh}H.$ (8)
Dataset Name Train/Test No. of Samples Generalization Evaluation
FaceForensics++ [30] Train 114884 -
Celeb-DF v1 (CDFv1) [23] Test 3136 Cross-domain
Celeb-DF v2 (CDFv2) [23] Test 16420 Cross-domain
FaceShifter (Fsh) [9] Test 8958 Cross-domain
Diffusion Models (DDPM, DDIM, LDM) Test 50,000 Novel Face Deepfake
Table 1: Dataset Information: The table provides a description of the datasets employed in our study to evaluate Cross-Domain and Unseen Deepfake Generalization.
Wavelet Classifier
$fv_{\text{low}}^{(n)},fv_{\text{high}}^{(n)}$ $=\text{DWT}(Z^{(n)})$ (9)
${fv^{\prime}}_{\text{low}}^{(n)}$ $=\text{MLP}(fv_{\text{low}}^{(n)})$ (10)
(a) ROC - DDIM
(b) ROC - DDPM
(c) ROC - LDM
Figure 2: AUCROC Plots: Receiver Operating Characteristic (ROC) curves for a) DDIM, b) DDPM, and c) LDM, depicting the models’ performance in terms of the Area Under the Curve (AUC), along with their true positive and false positive rates.
4 Set Up
In this section, we will first detail the evaluation protocol, evaluation metrics, and datasets used.
Dataset and Evaluation
In alignment with the training and evaluation protocol established by Yan et al. [33], the models undergo initial training on the FaceForensics++ c23 dataset [30]. Subsequent evaluations employ a cross-domain test using datasets such as Celeb-DF v1 (CDFv1) [23], Celeb-DF v2 (CDFv2) [23], and FaceShifter (Fsh) [9], thereby providing a robust framework to test the generalization capabilities of all the models. While the existing benchmarks do not encompass tests on emerging diffusion models, our research extends to examining generalizability with novel synthetic samples. Utilizing state-of-the-art diffusion models like DDPM [20], DDIM [14], and LDM [13] (wihtout text-guidance), we generate approximately 50,000 images from the CelebA dataset weights–none of which were included in training phase. This approach enables a comprehensive evaluation of model’s adaptability to novel and unseen data, leveraging its potential for practical deployment in digital forensics (Refer to Table 1).
Metrics
To assess the effectiveness of the results we use AUC (Area Under the Curve) and EER (Equal Error Rate) as fundamental metrics [33]. The AUROC represents the degree of separability achieved by the model, indicating how well the model can distinguish between normal and anomalous images. The AUROC is calculated as the area under the ROC curve, which plots the true positive rate (TPR) against the false positive rate (FPR) at various threshold settings [18]. The formula for AUROC can be expressed as,
[…]
The Equal Error Rate (EER) is a metric used to evaluate the performance of a binary classification system, especially in biometric verification or detection tasks. It is the point at which the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. EER is defined as,
$\text{EER}=\text{FAR}(\tau^{*})=\text{FRR}(\tau^{*}),$ (13)
where $\tau^{*}$ is the decision threshold that minimizes the difference between FAR and FRR. The EER provides a single scalar value to compare systems: lower EER indicates better system performance. It is often visualized on a Receiver Operating Characteristic (ROC) curve as the point where the curve intersects the line $\text{FPR}=1-\text{TPR}$.
@@ 4 Set Up
CLIP [11] CVPR-23 ViT [7] Self-Supervised 0.743 0.750 0.730 0.741
@@
Baselines
This study utilizes the state-of-the-art methods as baselines detailed by Yan et al. [33] for evaluating the performance of Wavelet-CLIP. These methods represent the current standard approaches for deepfake detection and are widely recognized in the research community. Broadly, the methods can be categorized into three types: naïve detectors, which employ traditional convolutional neural networks (CNNs) for direct classification; spatial detectors, which explore spatial artifacts or forgery regions in images; and frequency detectors, which analyze manipulation clues in the frequency domain. By using these standard methods—such as Xception [3], Capsule [26], F3Net [12]—as baseline models, I aim to ensure fair and transparent performance comparisons. Additionally, these baselines serve as a benchmark to assess both cross-domain generalization (using datasets like Celeb-DF v2 and FaceShifter) and unseen deepfake generalization (e.g., testing on large-scale datasets with 50k unseen samples). This approach ensures a robust comparison while highlighting improvements introduced by our Wavelet-CLIP.
5 Results
In this section, we will discuss the performance of our approach with various state-of-the-art approaches.
Cross-Data Performance
Table 2 clearly demonstrates the superior performance of Wavelet-CLIP compared to existing models. Among the supervised models, SRM [10] achieves the highest performance on Celeb-DF v1 (0.792) and Celeb-DF v2 (0.755), while other Xception-based methods, such as CORE [27] and UCF [17], exhibit competitive but slightly lower results. Notably, methods like MesoNet [1] and FFD [5] show limited generalization ability, achieving AUCs in the range of 0.609–0.711 on the datasets. Despite being trained end-to-end, these models rely heavily on backbone architectures like Xception [3] and ResNet50 [19], which may struggle to generalize under cross-domain settings due to their reliance on supervised learning and dataset-specific artifacts. The self-supervised models CLIP and Wavelet-CLIP demonstrate superior cross-dataset performance, indicating their robustness to unseen data. Unlike the supervised approaches, which require extensive training on specific datasets, self-supervised methods leverage pre-training on large-scale data, enabling better generalization to diverse domains.
@@ 5 Results
Specifically, for the CDFv1 dataset, traditional CLIP features fall short in capturing detailed representations; a standard ViT-L model [11] achieves an AUC of 0.743, while our model shows a significant improvement with a $+1.3\%$ increase. In every other scenario, our Wavelet-CLIP model stands out by consistently delivering strong performance. Notably, transformer-based models, including ours and those developed by Ojha [11], demonstrate effective representation capturing abilities for the FaceShifter (Fsh) dataset [9]. Interestingly, the best non-transformer model, Face X-ray [21], achieves an AUC of 0.655, whereas our approach exhibits a significant improvement with a $+7.7\%$ increase in AUC. The Fsh dataset, known for its sophisticated face manipulation techniques across diverse scenarios, presents a substantial challenge; yet, it appears that pre-trained transformers are particularly adept at discerning the subtle forgeries inherent in such deepfakes.
Robustness to Unseen Deepfakes
Next, Table 3 assesses the performance of Wavelet-CLIP on face images generated by unseen diffusion-based models. Among the supervised models, CapsuleNet [26] and SRM [10] stand out as strong performers, achieving average AUC scores of 0.768 and 0.651, respectively, with lower EER values compared to other supervised methods like MesoNet [1] and FFD [5]. However, methods such as Core [27] and F3-Net [12] exhibit significantly lower performance, with AUC values below 0.6, indicating their poor generalization capability when faced with unseen deepfake types. The CLIP model, which employs a self-supervised learning approach and uses the Vision Transformer (ViT) backbone, significantly outperforms the supervised methods. CLIP achieves an average AUC of 0.845 and a low EER of 0.235, demonstrating its robustness across all three datasets. This highlights the advantage of self-supervised learning in enhancing generalization to unseen data, particularly when compared to traditional supervised methods. Specifically, Wavelet-CLIP achieves the best performance on Celeb-DF v2 (0.759) and FaceShifter (0.732), and competitive results on Celeb-DF v1 (0.756). Compared to CLIP [11], which also employs a ViT-based backbone, Wavelet-CLIP consistently achieves higher AUC scores by leveraging wavelet-based features to capture both spatial and frequency domain artifacts.
@@
The aggregated performance of Wavelet-CLIP outperforms standard CLIP [11], showing an $+1.3\%$ increase in AUC and a $-1.2\%$ reduction in EER respectively. This highlights the substantial impact of integrating wavelet transformations within the classification head. Additionally, among the non-transformer based models, Capsule [26]—noted for its rotational invariance capabilities—performs best, yet it still falls short of matching Wavelet-CLIP, with performance differences of $12.5\%$ in AUC and $10.5\%$ in EER respectively. This further underscores the superior effectiveness of Wavelet-CLIP in handling sophisticated generative challenges. The incorporation of wavelet transforms enhances the model’s ability to detect subtle manipulations in images, such as high-frequency discrepancies often introduced by deepfake methods.
6 Discussion
The proposed Wavelet-CLIP framework introduces a novel approach to deepfake detection, combining wavelet-based frequency analysis with features derived from the Vision Transformer (ViT-L/14) pre-trained in a CLIP fashion [28]. The extensive experimental evaluations demonstrate that this integration significantly enhances the generalizability and robustness of the detection model, particularly in cross-domain and unseen deepfake scenarios. Unlike traditional supervised methods that rely on dataset-specific artifacts, Wavelet-CLIP leverages self-supervised representations to generalize across diverse generative families, setting a new benchmark for deepfake detection.
The Wavelet-CLIP is a task-agnostic feature extraction enabled by the CLIP-ViT encoder. Pre-trained on large-scale internet data, the frozen encoder provides strong transferable representations, allowing Wavelet-CLIP to excel across various datasets without the need for task-specific fine-tuning. Additionally, the integration of wavelet transforms enables the model to capture fine-grained frequency domain details, complementing spatial features and addressing a key limitation of previous methods like F3-Net [12]. However, the inclusion of wavelet decomposition and reconstruction introduces some computational overhead, which may limit real-time deployment in latency-sensitive applications. Moreover, while CLIP-derived features offer broad generalization, their performance might still depend on the diversity of the pre-training data, potentially limiting their applicability to niche or highly specialized deepfake artifacts. As observed, there is a noticeable improvement in AUC scores, and the model consistently achieves a significant edge in performance for all unseen deepfakes (Refer to Figure 2). In Figure 2, for certain cases, some models fall below the guessing threshold (AUC = 0.5) on the AUC curve, highlighting their limitations. In contrast, our approach demonstrates a clear advantage, with frozen CLIP-based features already showing a significant edge in detecting unseen facial deepfakes. However, Wavelet-CLIP further outperforms the existing state-of-the-art, establishing itself as a standalone solution.
Our current approach is specifically designed for detecting facial deepfake images, and it does not yet address other complex modalities such as audio-based, video-based, or audio-visual deepfakes, which remain significant challenges in the domain. Expanding our method to handle these modalities is crucial, as they often exhibit multi-modal inconsistencies that can be exploited for detection. Additionally, while our approach focuses on fine-tuning the wavelet-based classification head using pre-trained CLIP features, it does not leverage large-scale training on millions of fake images. Such large-scale training could further enhance the model’s ability to capture intricate forged features and improve detection performance. Consequently, our current work is limited in this regard. As part of future research, we identify these limitations as opportunities and stepping stones toward building more comprehensive, multi-modal, and robust deepfake detection systems capable of tackling emerging and diverse manipulation techniques. As a future work, we see that, incorporation of multimodal cues for deepfake detection could leverage models performance [29, 34]. Current detection frameworks rely exclusively on visual features; however, deepfakes often introduce inconsistencies in other modalities, such as audio, speech patterns, or facial expressions in video. For example, combining audio embeddings with visual representations could help detect deepfake videos where audio lip-sync mismatches are present. A multimodal CLIP approach, integrating both visual and auditory signals within a unified feature space, could represent the next step in building robust detection systems.